The market price in a perfect competetive industry is P =RM 10. Suppose that the total cost function of an individual firm in the industry is given by: TC= 100+ 5Q+ 0.02Q²

a) at what price and output would the firm maximize its profit?
b) what is the firm total profit at the profit maximizing price?
1
Expert's answer
2020-07-15T10:01:59-0400

a) at what price and output would the firm maximize its profit?


A firm in a perfect competitive firm will maximize their revenue when:

"\text{MR=MC}"

The marginal revenue of the firm will be as follows:

"\blue{\text{Revenue=Price * Quantity}}"

"\text{Revenue}=10Q"

The marginal revenue will therefore be:

"\text{Marginal revenue}=\dfrac{\delta\text{Revenue}}{\delta Q}=10"


The marginal cost on the other hand will be:

"\text{Marginal cost}=\dfrac{TC}{Q}=5+0.04Q"


Equating the MR=MC

"10=5+0.04Q"


"Q=\dfrac{10-5}{0.04}=125"


The profit maximizing quantity is 125.

The profit maximizing price will be equal to $10.

b) what is the firm total profit at the profit maximizing price?

The total profit will be equal to:

"\text{Profit function}=10-5-0.04Q"

"\text{Profit function}=10-5-0.04\times 125"

"\text{Profit function}=0"

The total profit at the price of 10 will be equal to 0.